# Semantic-Aware Power Allocation for Generative Semantic Communications with Foundation Models

### v1

@@ VI Conclusion
A generative SemCom framework for image tasks was proposed in this work, leveraging pre-trained foundation models for both semantic encoder and decode. Given the semantic encoder and decoder, the transmission reliability emerged as the primary factor influencing the perceptual quality of the regenerated images. Their mathematical relationship was modeled as a perception-error function, and the semantic values of the semantic data streams were defined. The perception-error function and the semantic values were empirically derived through numerical simulations on the Kodak dataset, providing a quantitative basis for further analysis. We investigated the semantic-aware power allocation problems and proposed two semantic-aware proportional and bisection methods. Numerical results demonstrated that the proposed semantic-aware bisection method consistently outperformed the semantic-aware proportional method and the conventional approach that treats the data streams equally with the same SNR. The performance advantages of the proposed semantic method become more pronounced with the increase of modulation order.

### v2

@@ VI Conclusion
A generative SemCom framework for image tasks was proposed in this work, leveraging pre-trained foundation models for both semantic encoding and decoding. Given the semantic encoder and decoder, the transmission reliability emerged as the primary factor influencing the perceptual quality of the regenerated images. Their mathematical relationship was modeled as a perception-error function, and the semantic values of the semantic data streams were defined accordingly. Both the perception-error function and semantic values were empirically derived through numerical simulations on the Kodak dataset, providing a quantitative basis for further optimization. We investigated the semantic-aware power allocation problem and proposed semantic-aware proportional and semantic-aware bisection methods. Numerical results demonstrated that the proposed semantic-aware bisection method consistently outperformed the semantic-aware proportional method and the conventional approach. The performance advantages of the proposed semantic method become more pronounced with the increase of modulation order.
VII Acknowledgement
This work was supported by the U.K. Department for Science, Innovation, and Technology under Project TUDOR (Towards Ubiquitous 3D Open Resilient Network).